OpenStudy (anonymous):

Each of these sequence or series formulas involves four quantities, for example, a1, an, d, r, n, and Sn. For each formula, describe the four quantities. Then explain how you can find the fourth quantity if you know the values of the other three. A) an=a1 + (n-1) d B) sn= n/2 ( a1-an) C) an=a1r ^ n-1 D) Sn = a1 ( 1-r^n) / 1-r Can anyone help me

OpenStudy (ikram002p):

what do u think ?

OpenStudy (ikram002p):

an=a1 + (n-1) d like this one :- 0,2,4,6,8,10.... as u can see here \(a_1=0\) and d =0 and u can get d like this :- 2-0=2 4-2=2 6-4=2 ... so according to this example what is \(n , a_0 ,d\) ?

OpenStudy (ikram002p):

typo :P \(a_1\)

OpenStudy (anonymous):

But how do I describe the 4 quantities? My teacher helped me get started with this but I'm still confused You solve for n by isolating it on the left side.... an = a1+nd-d an-nd = a1-d n(a-d) = a1-d n = (a1-d)/(a-d)